hamidiso23:

Complete the equation of the line through (2,-2) and (4,1) Use exact numbers. y = ?? solve for y I have to get it in slope intercept form y = mx + b

hamidiso23:

do I stat with y2-y1

hamidiso23:

x2 - x1

hamidiso23:

start

hamidiso23:

and find the slope then put it back in the equation?

hamidiso23:

@thisgirlpretty

ThisGirlPretty:

@umm

hamidiso23:

anyone?

ThisGirlPretty:

@dude

AN0NYM0US:

The incline catch type of the situation of a line is : y = mx + b The incline is addressed by m and the y-catch is addressed by b. You can utilize the directions of the two focuses to discover the incline: m = (y 2 - y 1) (x 2 - x 1) For this issue: m = (4 - 2 ) = 2 (1 - 2) - 1 To decide the y-block , substitute the directions of one or the other point into the slant capture condition utilizing an incline of - 2: y = mx + b y = - 2 • x + b Utilizing (1,4): 4 = - 2 • 1 + b 4 = - 2 + b b = 6 The condition of the line will be : y = - 2•x + 6 The incline catch type of the situation of a line is : y = mx + b The incline is addressed by m and the y-catch is addressed by b. You can utilize the directions of the two focuses to discover the incline: m = (y 2 - y 1) (x 2 - x 1) For this issue: m = (4 - 2 ) = 2 (1 - 2) - 1 To decide the y-block , substitute the directions of one or the other point into the slant capture condition utilizing an incline of - 2: y = mx + b y = - 2 • x + b Utilizing (1,4): 4 = - 2 • 1 + b 4 = - 2 + b b = 6 The condition of the line will be : y = - 2•x + 6

hamidiso23:

thats not the answer

hamidiso23:

-2 times x

hamidiso23:

u sure

hamidiso23:

that was wrong

hamidiso23:

it isnt -2x +6

AN0NYM0US:

X=-y-6/2

AN0NYM0US:

X=3 and y=6

hamidiso23:

thats wrong too

Serenity1jacksonsl:

@hamidiso23 wrote:
thats wrong too
he are she put the numbers wrong

hamidiso23:

@serenity1jacksonsl wrote:
@hamidiso23 wrote:
thats wrong too
he are she put the numbers wrong
who

AN0NYM0US:

Me

Serenity1jacksonsl:

i was looking at the top my bad

hamidiso23:

@serenity1jacksonsl wrote:
i was looking at the top my bad
u good

KjSaif:

do you remember the rise run method I told you ?

KjSaif:

first plot the lines on a graph

hamidiso23:

down or up and over right or left

hamidiso23:

there is no graph

KjSaif:

you can always make one

@hamidiso23 wrote:
there is no graph
you can always make one

IhelpUuHELPme:

Desmos

KjSaif:

^

hamidiso23:

ok

hamidiso23:

its not plotting

dude:

@hamidiso23 wrote:
do I stat with y2-y1
Yes \(m=\dfrac{y_2-y_1}{x_2-x_1}\) (2,-2) is point 1 and (4,1) is point 2

hamidiso23:

3/2

hamidiso23:

is the slope

dude:

Yes Now we substitute it into the general equation y=\(\dfrac32x+b\) From here you can choose either point 1 or 2 and substitute it in for x and y

hamidiso23:

and 4 is the y

KjSaif:

@hamidiso23 wrote:
its not plotting
No once you plot you can start to find your answer. You find the slope from there and then on.

dude:

If you choose point (4,1), 4 isn't the y

hamidiso23:

I choose

hamidiso23:

point 1

hamidiso23:

-2 + 2 equals zero

hamidiso23:

2+ 2 =4

hamidiso23:

isnt it 4 for the y intercept

dude:

Not quite, you just want to substitute it in the equation we got earlier \(y=\dfrac32x+b\) [From point 1, (2,-2)] 2 = x -2 = y You'll get: \(-2=\dfrac32\cdot2+b\) Does that make sense?

hamidiso23:

@dude wrote:
Not quite, you just want to substitute it in the equation we got earlier \(y=\dfrac32x+b\) [From point 1, (2,-2)] 2 = x -2 = y You'll get: \(-2=\dfrac32\cdot2+b\) Does that make sense?
so whats the answer

dude:

You want to isolate for b first

hamidiso23:

my brain hurts

dude:

xd What is \(\dfrac32\cdot2\)?

hamidiso23:

3

dude:

Yes So we have \(-2=3+b\) You subtract 3 from both sides

hamidiso23:

5

hamidiso23:

-5

dude:

Right So the equation of a line is y=mx+b We only need m and b (which we just solved) So we have \(y=\dfrac32x-5\)

hamidiso23:

tnx

hamidiso23:

tnx

hamidiso23:

Complete the equation of the line through (-6,-5)( -4, -4) Use exact numbers. y =?? I need to answer the question in y= mx+ b format

hamidiso23:

1/2 would be slope right

dude:

Yes

dude:

Now choose one of the points and substitute it in the general equation

hamidiso23:

point 1

hamidiso23:

( -6, -5)

dude:

\(-5=\dfrac12\cdot-6+b\) Do you know how to solve for b?

hamidiso23:

1/2

hamidiso23:

times -6?

dude:

Yes

hamidiso23:

-3

dude:

\(-5=-3+b\) Last step?

hamidiso23:

-8

dude:

Close You would have to add 3 on both sides

hamidiso23:

8

dude:

\(-5+3\)

hamidiso23:

-2

dude:

Righto Putting it back into the general equation \(y=\dfrac12x-2\)

hamidiso23:

tnx

hamidiso23:

Complete the equation of the line through (1,4) and (2,2)

hamidiso23:

Use exact numbers.

dude:

Lets make a new post

hamidiso23:

-2/1

hamidiso23:

is the slope right